Two 33ohms resistor are connected in parallel followed by two more 33ohms resistor connected in parallel. what value of a single resistor will be used to replace this 4 resistor

1
Expert's answer
2021-11-12T10:02:41-0500

Effective resistance "R_{eff}" of two resistors "R" in parallel is given by


"R_{eff}=\frac{RR}{R+R}=\frac{R^2}{2R}=\frac{R}{2}"


"R_{eff(1)}=\frac{33}{2}=16.5\Omega"


"R_{eff(2)}=\frac{33}{2}=16.5\Omega"


The value of single resistor

"=16.5+16.5"

"=33\Omega"